The upper boundaries for a distribution of waiting times (in seconds) in a grocery store aisle are 48, 61, 74, and 87. List the value for each lower boundary in this distribution
Answer #1

Upper boundries are 48, 61, 74, and 87
Class width = 61-48-1= 13-1=12

So lower boundries are
48-12 =36
61-12 = 49
74-12 = 62
87-12 = 75

So lower boundaries are 36,49,62,75

Please rate
Answer #2

in a grocery store aisle are 506274, and 86. List the value for each lower boundary in this distribution.